QUESTION 7 Assurme that the market for Good X is defined as follows: QD 64-16P and QS 16P-8. If the government imposes a price floor in this market at $3.00, what will consumer surplus be? $24.50 o $49.00 $9.00 $8.00 $32.00
Answer #1

Answer :Option D:$8

For price P=$3

Demand Qd=64-16*3=16 units

So

For consumer surplus = area of triangle abc=1/2*(4-3)*16=$8